A number added to twice the absolute value of the number is equal to 6. Find the
number.

Answer by Theo(13342)

x + 2 * |x| = 6

If x is greater than or equal to 0, then this formula becomes:

x + 2 * x = 6

This comes out to 3*x = 6 which makes x = 2

If x is less than 0, then this formula becomes:

x + 2 * (-x) = 6 which becomes x - 2 * x = 6

this comes out to be -x = 6 which comes out to be x = -6

You have two answers looks like.

x = 2 and x = -6

Let's see if they work.

Your original equation is x + 2 * |x| = 6

Let x = 2 and this equation becomes 2 + 2 * 2 = 6 which is true so the first answer is good.

Let x = -6 and this equation becomes -6 + 2 * 6 = 6 which becomes -6 + 12 = 6 which becomes 6 = 6 which is true so the second answer is good.

I would say that your answer is:

x = 2 or x = -6